OpenStudy (anonymous):

Which of the following is true? three hundred two degrees Fahrenheit is less than one hundred fifty degrees Celsius three hundred two degrees Fahrenheit is greater than one hundred fifty degrees Celsius three hundred two degrees Fahrenheit is equal to one hundred fifty degrees Celsius three hundred two degrees Celsius is less than one hundred fifty degrees Fahrenheit

OpenStudy (anonymous):

OK lets get the formulas :)

OpenStudy (anonymous):

For Fahrenheit we use: F = 9/5 C + 32 For Celsius we use: C = 5/9 (F - 32) Can you solve these now?

OpenStudy (anonymous):

@trenical ARE YOU THERE????

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Can you solve these problems by plugging in the numbers with the formulas I gave you?